OpenStudy (anonymous):

The landing of a spacecraft is cushioned with the help of airbags. During its landing on Mars, the velocity of downward fall is 16 meters/second. Immediately after the impact, the velocity is reduced to 1.2 meters/second. If the spacecraft has a mass of 11.5 × 10^4 kilograms, what is your estimate for the impulse if the time of impact is 0.8 seconds?

OpenStudy (souvik):

change in momentum = impulse on that object.

OpenStudy (anonymous):

but i dont know the force?

OpenStudy (souvik):

but you know the change in momentum...

OpenStudy (souvik):

F=m(v-u)/t or F*t=mv-mu impulse = change in momentum v=final velocity u= initial velocity